OpenStudy (anonymous):

Ryan left the science museum and drove south. Gabriella left three hours later driving 42 km/h faster in an effort to catch up to him. After two hours Gabriella finally caught up. Find Ryan's average speed.

OpenStudy (ragingsquirrel):

Gabriella went south, right?

OpenStudy (anonymous):

yeah

OpenStudy (ragingsquirrel):

It may be one of those dumb riddles. Ok lemme work on that

OpenStudy (anonymous):

They both will have travelled the same distance... Therefore Speed(Gabriella) * Time(Gabriella) = Speed(Ryan) * Time(Ryan) Speed(Ryan) = x km/h Time(Ryan) = 5 hours Speed(Gabriella) = (x+42) km/h Time(Gabriella) = 2 hours x*5 = (x+42)*2 5x = 2x + 84 3x = 84 x = 84/3 x = 28 km/h

OpenStudy (anonymous):

wow man thank you